### v3

# Explain the Black Box for the Sake of Science: the Scientific Method in the Era of Generative Artificial Intelligence

Abstract
The scientific method is the cornerstone of human progress across all branches of the natural and applied sciences, from understanding the human body to explaining how the universe works. The scientific method is based on identifying systematic rules or principles that describe the phenomenon of interest in a reproducible way that can be validated through experimental evidence. In the era of artificial intelligence (AI), there are discussions on how AI systems may discover new knowledge. We argue that human complex reasoning for scientific discovery remains of vital importance, at least before the advent of artificial general intelligence. Yet, AI can be leveraged for scientific discovery via explainable AI. More specifically, knowing what data AI systems deemed important to make decisions can be a point of contact with domain experts and scientists, that can lead to divergent or convergent views on a given scientific problem. Divergent views may spark further scientific investigations leading to new scientific knowledge.
keywords:
Explainable artificial intelligence, interpretability, scientific method, knowledge discovery
Human progress is inherently related to our ability to observe the world around us, and make sense of it. This frequently means identifying a set of rules or principles that describe systematically the phenomenon we are interested in. The word ‘systematic’ is crucial here: it means that the rules identified generalize to observations of the phenomenon that were unseen when deriving those rules. Indeed, this is the cornerstone of the modern scientific method galileo , and it commonly relies on mathematics, the ‘language’ that enables us to write systematic rules of various phenomena. The scientific method allows us to discover new knowledge about the world around us, from the natural sciences, including physics and biology, to the applied sciences, including medicine and engineering.
The scientific method is inherently connected to humans, as we have been the key players in discovering new principles (or theories) that explain the real world. Indeed, small modifications in our brain size and neuronal connections, made us the dominant species on Earth herculano2012remarkable . However, with the advent of and recent advances in artificial intelligence (AI), machines are also learning ‘principles’ (intended in a colloquial meaning) on scientific problems that underlie natural phenomena – see e.g.,  lam2023learning ; karagiorgi2022machine ; jumper2021highly ; schutt2019unifying ; esteva2019guide ; novakovsky2023obtaining ; merchant2023scaling ; yang2023large . Today, for a machine to learn ‘principles’ means to identify patterns in data and learn relationships that may generalize to the problem the machine is set to solve. The identification of these principles (i.e., patterns and relationships) is usually achieved via defining a learning task for an AI model, training the model on that task, and verifying that the trained model generalizes to unseen observations. Once this step is achieved, we are commonly interested in the results that the AI model gives us on the provided learning task, rather than on the ‘principles’ it used to reach those results. This is frequently satisfactory, and may be indeed our final goal. Yet, driven by curiosity, regulatory requirements, or else, we may want to know what ‘principles’ the machine used to obtain a certain result. To this end, the questions ‘what’ and ‘why’ are central to our discussion: (i) ‘what’ did the machine use to reach a certain decision, and (ii) ‘why’ did the machine use the ‘what’? The ‘what’ can be the input-output relationships identified by the machine, or just the data deemed important by the machine to reach a certain conclusion (when those explicit relationships are not available). In both cases, data is central. The ‘why’ is left to human complex reasoning to figure out, trying to interpret or explain the ‘what’. We argue that, for answering the ‘why’ question, and uncover the principles learned by the machine, data is the key, and it represents the point of contact between machines and humans, at least for the natural and applied sciences. We name the data the machine deemed important to reach a certain result, the machine view, as depicted in Fig. 1. The scrutiny of the machine view by human scientists and their complex reasoning can be the key to discovering new scientific knowledge.
To support this argument, we will dive into the scientific method that constitute the foundation of scientific discovery, and show how this can be revisited in the era of AI through explainability. We will further present a thought yet practically reproducible experiment that shows how this could be achieved by following some key requirements that we introduce.
We should also add that we share the concerns regarding the risks of adopting AI for scientific research (e.g., krenn2022scientific ; wang2023scientific ) recently brought forward by Messeri and Crockett messeri2024artificial , where they state that: “adopting AI in scientific research can bind to our cognitive limitations and impede scientific understanding despite promising to improve it”. We believe that explainable AI (XAI) is a way of bridging the gap between human knowledge and machine’s usage of the data, and can guide human experts to enrich human knowledge rather than being detrimental to it.
The scientific method in the era of AI
AI is commonly seen as a black-box tool, whereby humans struggle to understand why the machine took a certain decision or provided a certain forecast. This prevents us to understand possible ‘principles’ AI may have learnt, that could be useful to humans to enrich their knowledge or assess whether to trust the AI results. A way of addressing this issue, at least partly, is via providing a human-feasible pathway to ‘explaining’ AI, that could be achieved via XAI.
Explaining AI is a multifaceted and complex subject, and may have different meanings depending on e.g., the explanation’s end-users or the application area. In this work, we specifically address the explainability of AI for applications within the natural and applied sciences domain, where the intended end-users of the explanations are human scientists and domain experts. We argue that, for this audience of end-users, explaining AI means understanding ‘what’ data the machine deemed important to reach a certain result and ‘why’. The ‘how’ the machine used the ‘what’ should also be used, if available rudin2019stop . Answering the ‘what’ (and when possible the ‘how’) and ‘why’ questions can unveil those sought-for ‘principles’ the machine might have learnt. In fact, answering the ‘why’ question should be under the responsibility of human scientists and domain experts (see also the famous Richard Feyman’s discussion on “why questions” feynman1983why ).
The rationale for this argument is as follows. The scientific method, that underlies both the natural and applied sciences, is based on inductive reasoning, i.e., the formulation of hypothetical rules, and empirical evidence, i.e., observations and experiments, that validate the hypothetical rules. The hypothetical rules may be formulated before or after empirical evidence, and they can in fact be amended and improved after new empirical evidence comes to light. This approach to the scientific method is closely related to the prevailing philosophical view of science provided by Popper in his work “The Logic of Scientific Discovery” popper2005logic , although alternative views exist – see e.g., Feyereband feyerabend2020against , among others.
Figure 1: Comparative view of the scientific method workflow used by humans for scientific investigation and discovery [left, in green] and the machine (AI) workflow [right, in red]. Both workflows may lead to understand given problems in the natural and applied sciences, possibly leading to discovering systematic principles (or rules). We argue that comparing the existing human body of knowledge (what we call the human view) with the data used by the machine to make decisions (what we call machine view) may lead to further scientific investigation due to divergent views between human and machine views, that in turn can yield scientific discovery. It may also yield assessing results and deem them trustworthy for critical applications in the applied sciences, such as medicine. The key to obtain this convergent and/or divergent views is via explainable AI [yellow box on the bottom left], and more specifically outcome interpretability.
To better understand in practice how this approach works, let us take an example from physics, namely the law of universal gravitation introduced by Sir Isaac Newton in Philosophiae Naturalis Principia Mathematica newton . This law defines the force between two objects at a certain distance $r$ from each other, and each with its own mass, say $m_{1}$ and $m_{2}$
$F=G\frac{m_{1}m_{2}}{r^{2}}$ (1)
where $G$ is the constant of gravitation, that was accurately estimated in the famous Cavendish experiment almost 100 years after the law of universal gravitation was introduced. Newton in his book newton states that the law of universal gravitation he formulated was obtained using inductive reasoning derived from empirical observations.
This relatively simple example shows how the hypothetical rule of one of the most important discoveries in physics was defined after empirical evidence. Yet empirical evidence was necessary after the formulation of the hypothetical rule to confirm it, and to identify the value of the constant of gravitation $G$.
Another relevant example is the theory of relativity, formed by both special einstein1905electrodynamics and general einstein1916foundation relativity. These were introduced by Albert Einstein, who himself defined them as principle theories. The latter are theories formed by a body of knowledge (i.e., prior knowledge) that can guide understanding and discovery within a certain discipline. Indeed, Einstein used existing theoretical findings and some empirical evidence to build special and general relativity.
This second example highlights another important point. In addition to empirical evidence, hypothetical rules are formed together with prior knowledge.
In other fields of natural and applied sciences, we can draw similar conclusions. For instance, in biology, Gregor Mendel formulated the underlying rules of Mendelian inheritance after performing hybridization experiments on thousands of pea plants, and observing what traits were passed on to later generations of pea plants mendel1970versuche . In medicine, Alexander Fleming discovered penicillin, observing by chance that the fungal contamination of a particular bacterial culture seemed to kill the bacteria fleming1929antibacterial . In engineering, Alessandro Volta invented the battery following another scientist, Luigi Galvani, findings that a circuit of two metals can cause a frog’s leg to move.
From these examples, it should be clear that new knowledge in science is driven by real-world observations and the existing body of knowledge of a certain phenomenon of interest. These observations along with the existing knowledge lead to formulating hypothetical rules, that are successively validated by further observations, or falsified popper2005logic . The workflow just presented constitutes the foundation of the scientific method, and it ultimately leads to discovering new knowledge in the form of principles or theories, as depicted in Fig. 1 (left green box).
Observations, for what AI is concerned, are data and, in some cases, labels or known outputs, that the machine can use to learn a certain task, and become better at it by experience (i.e., training) mitchell1990machine . By training on some data (and labels), the machine learns some rules that are hopefully applicable to unseen data; in other words, the machine may have learnt some ‘principles’ in the form of patterns in data and relationships that generalize to unseen observations. The ‘principles’ the machine has learnt may or may not be those empirically-validated hypothetical rules that constitutes today’s human knowledge on a certain scientific problem. This machine workflow is depicted on the right red box in Fig. 1. Frequently, we are satisfied with the results AI is providing us, without further investigating why the machine provided those results – i.e., what ‘principles’ the machine used to obtain those results. For instance, two recent applications of AI to scientific problems that made the headlines are within the context of weather and climate science – weather forecasting lam2023learning – and biology – protein folding jumper2021highly . In these two applications, data is available, and it is frequently big, and our understanding of certain processes is lacking (e.g., cloud microphysics and protein folding rules), although a conspicuous body of prior knowledge is present. AI in these two applications, seems to outperform, in some cases, established methods built with prior human knowledge. In other words, AI seems to have learnt some underlying rules or principles (that is: patterns and relationships in data), that lead to better performance than achieved with human understanding of these problems. Hence, we might want to understand what those ‘principles’ are, and possibly close existing human-knowledge gaps. However, with the traditional workflow shown in Fig. 1 (right red block only) we are unable to achieve this step.
We argue that the key to bridge the gap between human and machine understanding is through XAI, that in turn can provide what we name here the machine view on a certain scientific problem (yellow box on the bottom right of Fig. 1). The machine view corresponds to ‘what’ data the machine deemed important to obtain a certain result, and it gives us a glimpse of those principles that the machine has learnt [^1] A glimpse and not the full picture, because to really retrieve those principles, we should respond to the ‘how’ the machine used those data deemed important. To achieve this, we would need to have full access to the inner workings of the AI system, along with all the relationships among variables and parameters that constitute the model; route that is practically not viable today.. We can then compare the machine view with the existing body of knowledge available (what we name human view), and try to respond the ‘why’ question, that is: why the machine deemed those data important. This comparison may lead to convergent views or divergent views on the problem being solved. Convergent views can reassure domain experts that AI results are trustworthy (aspect particularly critical in the applied sciences, such as in medicine, where wrong decision may mean the life or death of a patient). Divergent views can spark further scientific investigation. More specifically, scientists and domain experts may look at the machine view, and take a divergent (yet machine viable) understanding on what is commonly accepted in a certain field, thus providing possible pathways to fill knowledge gaps, and discover new knowledge. Going back to the weather forecasting and protein folding examples introduced earlier, it may allow us to fill the knowledge gaps in e.g., cloud microphysics (not likely given the spatial resolution and data used in current models) and protein folding rules, by identifying what data the machine deemed important to provide state-of-the-art predictions.
The workflow presented in Fig. 1 can open the path towards scientific discovery via XAI and open the exciting new field of XAI for Science. The latter naming convention is in analogy with AI for science, and scientific machine learning, that usually embeds domain-specific knowledge into AI models to produce plausible domain-constrained results raissi2019physics . XAI for Science (also referred to as scientific XAI) may use both constrained and unconstrained AI models, where the data the machine used to take a certain decision is available to the users.
However, in order for scientific XAI to have a chance of being successful, it needs to satisfy certain key pillars or requirements, that are necessary but not sufficient to achieve the task.
Explaining AI in Science: what + why
The starting point of scientific XAI is related to what we named the machine view, that answers the ‘what’ question: ‘what’ is the data the machine deemed important to obtain a certain result?
However, before diving into explainability and the machine view, we shall clarify a crucial assumption.
Assumption. In this perspective, we assume that the AI model has learnt something useful (i.e., some rules or ‘principles’ in the form of patterns and relationships) about the real-world phenomenon of interest. We further assume that those principles are encoded within the machine view. This translates into having a causal AI model that can accurately accomplish a scientific task (e.g., predict the future weather evolution or a certain protein structure) similarly or better than state-of-the-art human-knowledge-driven models.
The assumption just made simply implies that there is a chance of learning new rules in the form of patterns and relationships, using an AI model. In the assumption, we refer to a causal model. We note that the concept of causality, in machine learning, and more generally in science, remains widely debated pearl2018book , and may take different connotations and viewpoints einstein1916foundation ; bohr1937causality ; hume2000treatise ; popper2005logic . Therefore, causality should be assessed by scientists and domain experts, as part of their answer to the ‘why’ question.
Having framed our problem with the above assumption, we can now return to the machine view, noting that a useful and detailed approach to scientific inference using interpretable machine learning and its implications is provided by Freiesleben et al., freiesleben2022scientific .
The machine view can be retrieved via different methods within the context of XAI, where the term explainable [^2] We note that explanation and interpretation are frequently used interchangeably in machine learning, and we retain this ambiguity in the present work. However, several authors differentiate them, choice that may be useful in certain contexts. For instance, Rudin et al. rudin2019stop , define explainability as the field that tries to explain a posteriori black-box models, and interpretability as the field that build non-black-box models. (the ‘X’ in XAI) may have different meanings depending on the community and context, leaving room for a certain degree of ambiguity in the literature. In the English language, explanation may have two meanings according to the Oxford dictionary: 1) a statement or account that makes something clear, or 2) a reason or justification given for an action or belief. Knowing what data the machine deemed important to come to a certain result may be related to both definitions or, in fact, to none of them. On the one hand, it may render the result more clear or more trustable to an expert user. On the other hand, it may be the justification for a certain result, that may or not make sense to an expert user. In fact, both definitions may apply simultaneously. Yet, it may also provide none of the above, keeping 1) something unclear, or 2) without a reason or justification. Which of the above statements applies depends on whom the explanation is directed to, or better, on who is trying to understand the results of AI and the associated machine view. Hence, the machine view, strictly speaking, is only part of what we would define an explanation in the English language. The data the machine deemed important to obtain a certain result (the ‘what’ question) should be explained by human scientists/domain experts to make them clear and/or to give a reason or justification for those results. In other words, human scientists and domain experts, through their existing knowledge and complex reasoning, should respond to the question ‘why’ the machine used the ‘what’ identified via a variety of XAI methods.
We argue that it is exactly this latter aspect that makes AI results explainable, at least in the context of the natural and applied sciences: the human understanding and explanation of what data the machine deemed important. Therefore, explanation is the answer to both the ‘what’ and ‘why’ questions that we introduced in this work. We shall note that the ‘how’ the machine used the data it was given is as important as the ‘what’ rudin2019stop , but the ‘how’ is frequently difficult to obtain, yet when available it shall also be used along the ‘what’ data the machine deemed important to obtain a certain result.
A human scientist/domain expert can try to compare the machine view to their understanding and prior knowledge of the problem being tackled (human view). This comparison can then lead to a convergent view aligned with their human view, or to a divergent view. The latter may spark further investigation that can potentially lead to new knowledge.
Figure 2: Explainability workflow for domain experts and scientists. The point of contact between the machine and human experts is through data, and in particular the data deemed important by the machine to take decisions, what we call machine view. To obtain the machine view, we can use both post-hoc interpretability methods, that need to undergo evaluation due to accuracy issues, and self-explainable methods. These explainability methods lead to relevance maps or concepts that human scientists and domain experts can consult. These relevance maps or concepts can then be used to understand if they are in agreement (convergent views) or not (divergent views) with the human body of knowledge about a specific problem. The machine view needs to satisfy the three pillars of explainable AI for science (and specifically referred to scientists and domain experts): accuracy, reproducibility, understandability. The machine view example related to genetics (second panel from the top on the left-most box) is re-adapted from novakovsky2023obtaining .
However, before looking at what data the machine deemed important to make a certain prediction, we should be careful of not being fooled by the machine view. To avoid this, we should look more closely at recent XAI research and what the tools available are. XAI is part of the broader umbrella of AI transparency that encompasses different aspects, such as disclosing whether AI is used in a certain application, understanding the inner workings of AI models, and understanding what data AI models are using to make decisions yeung_2020 . The latter, also called model outcome explainability (interpretability), is the one of interest in this perspective.
Outcome explainability aims to understand how an AI model used its input features to produce a given output. Several strategies have been proposed to achieve this task. The two most prominent ones are (i) methods to understand black-box models, usually referred to as post-hoc interpretability methods, and (ii) intrinsically- or self-explainable models (also referred to as white-box models). In Fig. 3, we show the XAI workflow relevant to this work, constituted by outcome explainability via post-hoc interpretability methods and/or self-explainable methods, that can give us the required machine view in the form of relevance maps or concepts.
Post-hoc interpretability methods typically provide a relevance or saliency map that reveals the importance of each input feature samek2021explaining to a certain output. Post-hoc interpretability methods have some desirable properties: they are model agnostic (e.g., DeepLIFT shrikumar2017learning , SHAP lundberg2017unified , GradCAM selvaraju2017grad , Integrated Gradients sundararajan2017axiomatic , LIME ribeiro2016should , LRP bach2015pixel ), and they do not modify the underlying AI model, as they are applied a posteriori. However, they may provide unfaithful explanations of the model outcomes sanity_adebayo ; reliable_dylan , and they tend to be sensitive to negligible perturbations ghorbani2019interpretation ; dombrowski2019explanations . Therefore, the accuracy of their explanations needs to be carefully evaluated turbe2023evaluation . Also, their explanations may be non-unique rudin2019stop , further demanding for ‘how’ data was used, vs just ‘what’ data was deemed important.
Self-explainable models provide concepts that are understandable by humans, and provide explanations that are more faithful than post-hoc interpretability methods rudin2019stop . Some examples of these methods include: decision trees and rule-based models (e.g., letham2015interpretable ; zeng2017interpretable ,) generalized additive models (e.g.,agarwal2021neural ), and prototype-based methods (e.g., nauta2023pip ; turbe2024protos ). However, these models may not perform as well as black-box models (e.g., deep neural network architectures) for certain complex scientific tasks (e.g., lam2023learning ; jumper2021highly ), thereby limiting their applicability, although this point of view is a source of lively debate rudin2019stop . For further reading on self-explainable methods and post-hoc interpretability the interested reader can refer to rudin2022interpretable .
The issues introduced are clearly undesirable for the objective we are trying to achieve here. For instance, they may lead to scientific hallucinations if we rely on inaccurate explanations (leading to some of the concerns brought forward by Messeri and Crockett messeri2024artificial ), or they may not be reproducible nor understandable, and give insights that are worse than our current understanding, thereby supplying useless information to domain experts and scientists. These key points, constitute the bases to define the foundational pillars that we require for scientific XAI to avoid being fooled, and they apply to both post-hoc interpretability methods and self-explainable models. Before specifying these pillars, we should point out that several authors proposed requirements that XAI should satisfy for a range of applications – the interested reader can consult e.g., phillips2021four ; rudin2022interpretable ; dwivedi2023explainable ; cambria2023seven , among others. In this perspective, we focus on the key pillars in the context of XAI for science, focusing specifically on the machine view, that conflates both model as well as explanation requirements.
Proposition. Requirements for the machine view within XAI for Science.
  1. 1.
    Accuracy. The machine view should be accurate. In both post-hoc interpretability and self-interpretable methods, practitioners tend to assume that the machine view is accurate, and we are able to correctly capture the data portions or concepts deemed important by the machine. This is frequently a wrong a assumption – for instance, post-hoc interpretability methods provide results that may be widely different depending on the method used. Similarly, self-interpretable methods may provide a set of explanations (or concepts) that is constrained by human prescribed instructions, thereby possibly limited.
  2. 2.
    Causality. The machine view should derive from a causal model, when this is appropriate, to avoid that the data deemed important by the machine is just related to spurious correlations. We stress that the causality requirement we set here should be part of the human scientist/domain expert assessment when attempting to answer the question ‘why’, as specified in our assumption on the model.
  3. 3.
    Reproducibility. The machine view should be reproducible, at least in a statistical sense, which means that if a certain outcome explanation is given for a certain learning task, this can be systematically reproduced. This requirement is important as it provide a robust machine view as opposed to one that can be easily changed, and it is an key building block of the scientific methods (including in the applied sciences, such as medicine, where statistical reproducibility is required).
  4. 4.
    Understandability. The machine view should be understandable to scientists and domain experts. Let me rephrase. If we want a scientist to make sense of the data used by a machine, this data should contain viable features that allow a scientist to tap into its existing corpus of knowledge. This requirement is a common desiderata in XAI, and it is frequently referred to as transparency. In the context of this perspective, we seek a distinction since we are specifically targeting scientists and domain experts, and their understanding of a given scientific problem.
The first requirement is not new; it is a basic requirement and starting point for XAI (let alone scientific XAI), and several frameworks have been deployed to evaluate interpretability results – see e.g., turbe2023evaluation among others. The other three requirements are instead more tailored towards XAI for science, and specify how to interface the machine view to domain experts and scientists. We note that causality is gaining traction, given its importance and implications for XAI – see for instance beckers2022causal ; carloni2023role . By having these pillars in place, scientists may grasp divergent views, increase their understanding of a given problem, and possibly fill knowledge gaps, provided that the assumption set out at the beginning of this section is satisfied.
We shall note that the complementary reasoning is also true. Scientists and domain experts may have a convergent view with the machine, thereby they may use the machine view as a tool for trustworthiness. The latter is a topic that is becoming increasingly important and regulated, especially for critical fields in the applied sciences, such as medicine eu2021aiact .
A practical example of XAI for science
So far, we introduced the perspective of why XAI may be a possible way forward in closing certain knowledge gaps in the natural and applied sciences, without showing any tangible example on how this can be achieved. In this section, we would like to introduce some experiments carried out in my group, where we used XAI, and post-hoc interpretability in particular, in the context of extreme weather prediction. We would like the reader to see this as a thought experiment, rather than entering into the details on how the results were obtained, noting that we provide an open-source repository with the details and code to fully reproduce the results.
The framing of the problem is as follows. We want our AI approach to predict heatwaves in the Indochina peninsula 1 day ahead of time, using the prior 5 days. The associated learning task for our AI framework (based of a transformer neural network) is to correctly classify whether 1 day in the future corresponds or not to the onset day of a heatwave. We train our AI framework on time series data corresponding to 4 different variables that are relevant to heatwaves for the 5 days prior to the onset (or no onset) of a heatwave, and on a geographical domain larger than the Indochina peninsula. The training dataset is composed of 365 samples, the validation set of 77 samples, and the testing set also of 77 samples. The classification results show an overall accuracy on the test set of 88.3% (68/77), with a true positive (heatwave event) rate of 76.5% (13/18) and a true negative (no event) rate of 91.6% (55/59). From these results, we may argue that the machine learnt some ‘principles’ (statistical patterns in the data used) related to heatwaves in Indochina, and satisfies the assumption that we set in the previous section. We also note that, a transformer neural network may provide a causal model nichani2024transformers , and many works are devoted to this topic – see e.g., melnychuk2022causal ; rohekar2024causal .
We would like now to understand what data the machine deemed important to learn these ‘principles’. To this end, we adopt post-hoc interpretability, and in particular the methods and evaluation framework presented in turbe2023evaluation , to guarantee that the relevance maps (i.e., the machine view) are correctly identified. We then compute the average machine view across all samples that are correctly classified, for the 4 variables considered, and for the 5 days prior to a heatwave striking that were used as inputs. We repeat the experiment a number of times, obtaining results that are similar to each other.
The first question we should ask ourselves is whether the ‘machine view’ just retrieved satisfies the pillars or requirements of XAI for science set out in the previous section. By evaluating the accuracy of the relevance maps, we attempt to satisfy the first requirement, namely accuracy. The possible inherent causality of the transformer model satisfies the second requirement, that we deem appropriate for this application. The reproducibility of the results for different repetitions of the experiment satisfies the third, namely reproducibility. Finally, the variables used are understandable to domain experts (i.e., weather and climate scientists), satisfying the fourth requirement. It is clear that further work should be undertaken to guarantee that the above requirements are indeed satisfied. But, for the sake of this thought experiment, we shall assume that this is the case.
What can we learn from the ‘machine view’ just retrieved? We can look at what domain experts understand of heatwaves in Indochina. To this end, we found only one relevant reference luo2018synoptic , that shows how temperature anomalies in South China seem to be systematic precursors of Indochina’s heatwaves, and constitutes the human view on this problem (as reported in Fig. 3). Remarkably, this human view and the machine view are convergent, at least partially – they deem the same temperature data as important for heatwaves in Indochina in the days preceding these events, as depicted in Fig. 3 (convergent views). We note that in the figure, for easier readability, we only show 3 days instead of 5. Interestingly, the machine also deem the water vapour as an important precursor to heatwaves in Indochina – more specifically, water vapor in India. Indeed, in this case the human view and the machine view are divergent, as depicted in Fig. 3 (divergent views). In the figure, we do not report the other two variables for the sake of readability.
Figure 3: Prediction of the onset of heatwaves in the Indochina Peninsula via a transformer neural network. We compare the associated machine view (i.e., the data the machine used to make its prediction) via post-hoc interpretability methods [bottom two yellow blocks] with the domain-expert view obtained from a scientific paper [top green box]. The machine view is convergent (i.e., in agreement) with domain scientists for temperature, while it is divergent (i.e., in disagreement) for water vapor.
A way of seeing the simple experiment presented here is the following. For temperature, (human) domain knowledge luo2018synoptic partially overlaps the machine view, and they seem to provide a convergent (albeit not fully) view on the problem. For water vapour, (human) domain knowledge luo2018synoptic does not overlap the machine view, and provide a divergent view on the problem (we shall note that we are working with domain experts on the specific example presented here).
The divergent view can promote further scientific investigation that may lead to new knowledge on the topic, and it is relevant for both the natural and applied sciences, especially in those fields and processes where we have limited knowledge. The convergent view may reassure us that the results are trustworthy, and it is particularly relevant for critical applications in the applied sciences, such as in medicine, where wrong decisions may have catastrophic consequences.
Obviously, the relatively simple example shown here has some limitations. For instance, the accuracy of the machine view may be further tested, with e.g., other methods, the causality of the model should be further explored, and variables as well as the data configurations used should be expanded, to further assess the reproducibility of the problem and improve its understandability. These limitations showcase the challenges that scientific XAI can encounter, and that needs to be tackled for this endeavour to have a chance of succeeding. We shall discuss this aspect next.
Limitations and future outlook
From the example just introduced, it is clear that scientific XAI has some limitations and possible pitfalls, especially given the current approaches to XAI. I shall discuss them in the following, using the four requirements set for scientific XAI as a roadmap, and providing a final comprehensive future outlook on the subject.
Let us start from the accuracy requirement. To obtain what we called the machine view there exist several methods, grouped into two different categories, namely post-hoc and intrinsic (also referred to as ante-hoc) interpretability – see e.g., graziani2023global for a comprehensive review and taxonomy of interpretable AI. Post-hoc interpretability suffers from inaccuracy, aspect that may render futile the effort by domain expert to explain why certain data was used, if the data pinpointed is wrong. To this end, accuracy evaluation methods are as important, see e.g., turbe2023evaluation . Intrinsic interpretability may address the accuracy issues that plague post-hoc interpretability, yet they may be model-specific, human-constrained, and unable to provide machine views for complex deployed models that may not be easily accessible. Yet, several promising works are being carried out in the context of intrinsic interpretability – see e.g., alvarez2018towards ; zhang2022protgnn ; turbe2024protos . Another recent promising area is (intrinsic) compositional interpretability tull2024towards , that may help tie together different intrinsic interpretability viewpoints under a unified framework.
In terms of the causality requirement, the patterns and relationships identified by the machine may be just spurious correlations, without any causality, and may fool scientists towards chasing wrong answers to the ‘why’ question. Understanding whether causality is appropriate to the problem being investigated, and assessing whether the machine view is causal is a key aspect that scientists should take into account in answering ‘why’ the machine deemed important certain data. Indeed, several works are now exploring the intersection between XAI and causality carloni2023role , two fields that in the computer science community have drifted apart over the years beckers2022causal , but that should be kept together, especially in the natural and applied sciences. To this end, some promising works are emerging – see e.g., scholkopf2022causality ; janzing2020feature ; xu2020causality . In addition, Peters et al. book peters2017elements provides an introduction to causality for learning algorithms, while Pearl’s book offers an accessible roadmap and perspective on causal inference from both a statistical and a philosophical perspective pearl2018book .
In terms of reproducibility, we should be able to retrieve the same machine view repeatedly, for a given task and set of samples. In this context, some issues derive from the possible non-uniqueness of the machine view that interpretability methods may provide (especially post-hoc interpretability methods). This can hinder our understanding of why those machine views were used. To this end, answering the question ‘how’ was the machine view (i.e., the data deemed important) used by the machine can provide a more useful insight, as well pointed out by Rudin rudin2019stop .
Finally, in terms of understandability, the machine view should allow scientists and domain expert to tap into their body of knowledge, to allow for comparing machine and human views on a certain phenomenon of interest. This is partly connected to what discussed by Freiesleben et al. freiesleben2022scientific . More specifically, the model should be able to address a question understandable to scientists, with a machine view that can also be understood by scientists within their framework of domain-specific knowledge. The choice of data and data configuration is therefore critical, and frequently arbitrary, leaving a degree of subjectivity and arbitrariness to the problem. For instance, if we consider the growing field of domain adaptability (i.e., using an AI model for a different domain than the one it was trained on) long2015learning , the machine view may not be understandable to scientists. These issues are less pronounced in the context of fine-tuning a certain model, as the pre-training is usually accomplished within the same domain as the fine-tuning, hence understandable to domain experts working in the specific field.
The challenges just presented, together with the overall premise of being able to transfer knowledge to human scientists via the machine view we introduced, may let scientific XAI seem like a futile endeavour. Yet, we believe that there is concrete hope. Let me take a brief detour, and explain why.
The perspective presented here is built on several works that have been published in the past few years within the context of AI, but perhaps one contribution stands out: the inspiring New York Times editorial by Steven Strogatz covering the winning of AlphaZero against Stockfish strogatz2018one . In that piece, Strogatz states: “What is frustrating about machine learning, however, is that the algorithms can’t articulate what they’re thinking. We don’t know why they work, so we don’t know if they can be trusted. AlphaZero gives every appearance of having discovered some important principles about chess, but it can’t share that understanding with us.”. He additionally cites Garry Kasparov (the former world chess champion) that stated: “we would say that its [AlphaZero] style reflects the truth. This superior understanding allowed it to outclass the world’s top traditional program despite calculating far fewer positions per second.” kasparov2018chess .
A few years later from that editorial, researchers showed that Go players are improving their game looking at how AlphaGo, DeepMind’s Go machine is playing, and providing them with a divergent view shin2023superhuman .
While science is a rather different application than Go, we believe that leveraging explanation for scientific inquiry that may ultimately lead to knowledge discovery may not be a foolish path. XAI is a possible path forward to keep complex human reasoning in the loop while bridging current AI power in discovering patterns and relationships in data. XAI may also alleviate some of the risks that we may face when using AI for scientific discovery, that we share with Messeri and Crockett messeri2024artificial .
We add that some views suggesting that current generative AI, including ChatGPT, and foundation models can autonomously generate new science may not be a plausible path unless complex human reasoning (provided by human scientists and domain experts) is kept in the loop, at least until when we will achieve artificial general intelligence Superintelligence .
Acknowledgments
We want to thank my two PhD students, Jiawen Wei and Chenyu Dong for many of the results presented in this work. We thank all members of my group at NUS, MathEXLab, for their dedication to research that made this work possible. We also thank all the colleagues that provided constructive criticism, from several different research fields and perspectives, that helped improve this manuscript significantly. We acknowledge funding from MOE Tier 2 grant 22-5191-A0001-0 ‘Prediction-to-Mitigation with Digital Twins of the Earth’s Weather’ and MOE Tier 1 grant 22-4900-A0001-0 ‘Discipline-Informed Neural Networks for Interpretable Time-Series Discovery’.