### v4

Abstract
The scientific method is the cornerstone of human progress across all branches of the natural and applied sciences, from understanding the human body to explaining how the universe works. The scientific method is based on identifying systematic rules or principles that describe the phenomenon of interest in a reproducible way that can be validated through experimental evidence. In the era of generative artificial intelligence, there are discussions on how AI systems may discover new knowledge. We argue that human complex reasoning for scientific discovery remains of vital importance, at least before the advent of artificial general intelligence. Yet, AI can be leveraged for scientific discovery via explainable AI. More specifically, knowing the ‘principles’ the AI systems used to make decisions can be a point of contact with domain experts and scientists, that can lead to divergent or convergent views on a given scientific problem. Divergent views may spark further scientific investigations leading to interpretability-guided explanations (IGEs), and possibly to new scientific knowledge. We define this field as Explainable AI for Science, where domain experts – potentially assisted by generative AI – formulate scientific hypotheses and explanations based on the interpretability of a predictive AI system.
keywords:
Explainable artificial intelligence, interpretability, scientific method, knowledge discovery
Human progress is inherently related to our ability to observe the world around us, and make sense of it. This frequently means identifying a set of rules or principles that describe systematically the phenomenon we are interested in. The word ‘systematic’ is crucial here: it means that the rules identified generalize to observations of the phenomenon that were unseen when deriving those rules. Indeed, this is the cornerstone of the modern scientific method galileo , that has allowed us to discover new knowledge in various fields, from the natural sciences, including physics and biology, to the applied sciences, including medicine and engineering.
The scientific method is inherently connected to humans, as we have been the key players in discovering new principles (or theories) that explain the real world. However, with the advent of and recent advances in artificial intelligence (AI), machines are also learning ‘principles’ (intended in a colloquial meaning) on scientific problems that underlie natural phenomena – see e.g.,  lam2023learning ; karagiorgi2022machine ; jumper2021highly ; schutt2019unifying ; esteva2019guide ; novakovsky2023obtaining ; merchant2023scaling ; yang2023large . Today, for a machine to learn ‘principles’ means to identify patterns in data and learn relationships that may generalize to the problem the machine is set to solve. The identification of these principles (i.e., patterns and relationships) is usually achieved via defining a learning task for an AI model, training the model on that task, and verifying that the trained model generalizes to unseen observations. Once this step is achieved, we are commonly interested in the results that the AI model gives us on the provided learning task, rather than on the ‘principles’ it used to reach those results. This is frequently satisfactory, and may be indeed our final goal. Yet, driven by curiosity, regulatory requirements, or else, we may want to know what ‘principles’ the machine used to obtain a certain result. To this end, the questions ‘what’ and ‘why’ are central to our discussion: (i) ‘what’ did the machine use to reach a certain decision, and (ii) ‘why’ did the machine use the ‘what’? The ‘what’ can be the input-output relationships identified by the machine, or just the data deemed important by the machine to reach a certain conclusion (when those explicit relationships are not available). In both cases, data is central. The ‘why’ is left to human complex reasoning to figure out, trying to interpret or explain the ‘what’. We argue that, for answering the ‘why’ question, and uncover the principles learned by the machine, data is the key, and it represents the point of contact between machines and humans, at least for the natural and applied sciences. We name the data the machine deemed important to reach a certain result, the machine view, as depicted in Fig. 1. The scrutiny of the machine view by human scientists and their complex reasoning can be the key to discovering new scientific knowledge. We shall make an important observation here: ‘how’ the machine used the information/data it is given is also an crucially important. However, that information is frequently not available to us, and deriving it may be quite challenging. Therefore, for this manuscript, we focus on the ‘what’ and ‘why’ questions, noting that the ‘how’ is equally important, when available.
To support our argument, we will dive into the scientific method that constitute the foundation of scientific discovery, and show how this can be revisited in the era of AI through explainability. We will further present a thought experiment (yet practically reproducible) that shows how this could be achieved, while highlighting some potential limitations.
We should also add that we share the concerns regarding the risks of adopting AI for scientific research (e.g., krenn2022scientific ; wang2023scientific ) recently brought forward by Messeri and Crockett messeri2024artificial , where they state that: “adopting AI in scientific research can bind to our cognitive limitations and impede scientific understanding despite promising to improve it”. We believe that explainable AI (XAI) is a way of bridging the gap between human knowledge and machine’s usage of the data, and can guide human experts to enrich human knowledge rather than being detrimental to it.
The scientific method in the era of AI
AI is commonly seen as a black-box tool, whereby humans struggle to understand why the machine took a certain decision or provided a certain forecast. This prevents us to understand possible ‘principles’ AI may have learnt, that could be useful to humans to enrich their knowledge or assess whether to trust the AI results. A way of addressing this issue, at least partly, is via XAI, where the latter is composed of two elements: (i) interpretability, and (ii) explanation of interpretability results in human-understandable terms; the two together provide a human-feasible pathway to ‘explaining’ AI, and yield interpretability-guided explanations (IEGs). The first element, interpretability, aims to answer the question ‘what’ (and possibly ‘how’, when available rudin2019stop ) data the machine deemed important to reach a certain result. The second element, explanation of interpretability results is left to domain experts and scientists that can generate hypotheses, if they have a divergent understanding of the data space compared to the machine. Answering the ‘what’ and ‘why’ questions can unveil those sought-for ‘principles’ the predictive AI model might have learnt.
The rationale for this argument is as follows. The scientific method, that underlies both the natural and applied sciences, is based on inductive reasoning, i.e., the formulation of hypothetical rules, and empirical evidence, i.e., observations and experiments, that validate the hypothetical rules. The hypothetical rules may be formulated before or after empirical evidence, and they can in fact be amended and improved after new empirical evidence comes to light. This approach to the scientific method is closely related to the prevailing philosophical view of science provided by Popper in his work “The Logic of Scientific Discovery” popper2005logic , although alternative views exist – see e.g., Feyereband feyerabend2020against , among others.
Figure 1: The scientific method in the era of AI. Comparative view of the approach used by humans (left, green) and AI (right, grey) to identify scientific principles that can generalize to unseen observations/data. Comparing the human view (existing knowledge) with the machine view (AI-used data) can drive new investigations from divergent perspectives or validate results for critical applications like medicine. The key to this process is interpretability (red box, bottom right), that can give us the machine view. Through this, we can anchor human explanations to the machine view, leading to interpretability-guided explanations (IGEs), that may lead to knowledge discovery.
To understand this approach in practice, consider Newton’s law of universal gravitation, which defines the force between two masses as $F=Gm_{1}m_{2}/r^{2}$. Newton derived this law inductively from empirical observations newton , where the gravitational constant $G$ was only determined later through Cavendish’s experiment.
Similarly, Einstein’s theories of special einstein1905electrodynamics and general relativity einstein1916foundation , which extended Newtonian mechanics, were developed using both theoretical insights and empirical evidence. These examples highlight how scientific discovery often combines hypothesis formation, prior knowledge, and empirical validation.
This pattern appears across disciplines. In biology, Mendel deduced inheritance laws through pea plant experiments mendel1970versuche . In medicine, Fleming discovered penicillin by observing bacterial inhibition due to fungal contamination fleming1929antibacterial . In engineering, Volta’s battery was inspired by Galvani’s findings on electrical stimulation in frogs.
These examples illustrate how scientific knowledge emerges from real-world observations and prior knowledge. Hypothetical rules are formulated, then validated or refuted through further evidence popper2005logic – a process that underpins the scientific method and drives discovery in the form of principles or theories (Fig. 1, left green box).
Observations, for what AI is concerned, are data and, in some cases, labels or known outputs, that the machine can use to learn a certain task, and become better at it by experience (i.e., training) mitchell1990machine . By training on some data (and labels), the machine learns some rules that are hopefully applicable to unseen data; in other words, the machine may have learnt some ‘principles’ in the form of patterns in data and relationships that generalize to unseen observations.
The ‘principles’ AI learns may or may not align with empirically validated scientific knowledge (Fig. 1, right red box). Often, we are satisfied with AI results without investigating why they were obtained – i.e., what principles the machine relied on.
Two notable AI applications in science – weather forecasting lam2023learning and protein folding jumper2021highly – showcase this. Both fields have vast data and theoretical understanding. Despite this, AI has in some cases outperformed traditional methods based on human knowledge, suggesting it has identified underlying patterns that improve performance. Hence, we might want to understand what those ‘principles’ are, and possibly close existing human-knowledge gaps. However, with the traditional workflow shown in Fig. 1 (right red block only) we are unable to achieve this step.
We argue that the key to bridge the gap between human and machine understanding is through interpretability methods applied to the AI system, that in turn can provide what we name here the machine view on a certain scientific problem (yellow box on the bottom right of Fig. 1). The machine view corresponds to ‘what’ data the machine deemed important to obtain a certain result, and it gives us a glimpse of those principles that the machine has learnt [^1] A glimpse and not the full picture, because to really retrieve those principles, we should respond to the ‘how’ the machine used those data deemed important. To achieve this, we would need to have full access to the inner workings of the AI system, along with all the relationships among variables and parameters that constitute the model; route that is practically not viable today.. We can then compare the machine view with the existing body of knowledge available (what we name human view), and try to respond the ‘why’ question, that is: why the machine deemed those data important. This comparison may lead to convergent views or divergent views on the problem being solved. Divergent views can spark further scientific investigation. More specifically, scientists and domain experts may look at the machine view, and take a divergent (yet machine viable) understanding on what is commonly accepted in a certain field, thus providing possible pathways to fill knowledge gaps, and discover new knowledge.
The workflow presented in Fig. 1 can open the path towards scientific discovery via XAI, also referred to as XAI for Science. The latter naming convention is in analogy with AI for science, and scientific machine learning, that usually embeds domain-specific knowledge into AI models to produce plausible domain-constrained results raissi2019physics . XAI for Science (also referred to as scientific XAI) may use both constrained and unconstrained AI models, where the data the machine used to take a certain decision is available to the users.
However, in order for scientific XAI to have a chance of being successful, it needs to satisfy certain key pillars or requirements, that are necessary but not sufficient to achieve the task.
Explainable AI for Science: what (and how) + why
The starting point of XAI for Science is related to what we named the machine view, that answers the ‘what’ question: ‘what’ is the data the machine deemed important to obtain a certain result? However, before diving into explainability and the machine view, we shall clarify a crucial assumption.
Assumption. In this work, we assume that the predictive AI model has learnt something useful (i.e., some rules or ‘principles’ in the form of patterns and relationships) about the real-world phenomenon of interest. We further assume that those principles are encoded within the machine view. This translates into having a causal AI model that can accurately accomplish a scientific task (e.g., predict the future weather evolution or a certain protein structure) similarly or better than state-of-the-art human-knowledge-driven models.
The assumption just made implies that there is a chance of learning new rules in the form of patterns and relationships, using an AI model. In the assumption, we refer to a causal model. We note that the concept of causality, in machine learning, and more generally in science, remains widely debated pearl2018book , and may take different connotations and viewpoints einstein1916foundation ; bohr1937causality ; hume2000treatise ; popper2005logic . Therefore, causality should be carefully assessed, by e.g., AI experts, as well as by scientists and domain experts, as part of their answer to the ‘why’ question.
Having framed our problem with the above assumption, we can now return to the machine view, and on how this can be retrieved. The relevant area of AI research for this task is called outcome interpretability, which explains how AI models use input features to generate outputs yeung_2020 . The two main approaches within this context are: (i) post-hoc interpretability and (ii) ante-hoc interpretability.
Post-hoc interpretability (e.g., DeepLIFT shrikumar2017learning , SHAP lundberg2017unified , GradCAM selvaraju2017grad ) generate saliency maps (i.e., machine views) to highlight important input features samek2021explaining . They are model-agnostic and do not alter AI models but may produce unreliable and non-unique explanations sanity_adebayo ; reliable_dylan ; rudin2019stop , requiring careful validation turbe2023evaluation .
Ante-hoc interpretability (e.g., decision trees letham2015interpretable , generalized additive models agarwal2021neural , prototype-based methods nauta2023pip ) may provide a more faithful machine view but may underperform in complex tasks lam2023learning ; jumper2021highly . This trade-off remains a topic of debate rudin2019stop . For further reading on these methods, see rudin2022interpretable .
The issues just outlined pose challenges for our overall objective. Relying on inaccurate explanations can lead to scientific hallucinations messeri2024artificial , while non-reproducible or unclear insights may provide little value to scientists.
To prevent such pitfalls, we define foundational pillars for scientific XAI, applicable to both post-hoc interpretability and self-explainable models. While various authors have proposed XAI requirements for different applications phillips2021four ; rudin2022interpretable ; dwivedi2023explainable ; cambria2023seven , we focus on three key pillars that the machine view should have for XAI for Science to succeed, accuracy, reproducibility, and understandability, or in brief ARU:
  1. 1.
    Accuracy. The machine view should be accurate. In both post-hoc and ante-hoc interpretability, practitioners tend to assume that the machine view is accurate; assumption that is frequently wrong.
  2. 2.
    Reproducibility. The machine view should be reproducible, at least in a statistical sense, which means that if a certain outcome explanation is given for a certain learning task, this can be systematically reproduced.
  3. 3.
    Understandability. The machine view must be understandable to scientists and domain experts, meaning it should present viable features that connect to existing knowledge. This aligns with the common XAI desideratum of transparency, but here, we emphasize its relevance to scientific understanding.
The first requirement is not new, and several frameworks have been deployed to evaluate interpretability results – see e.g., turbe2023evaluation among others. The other two requirements are instead more tailored towards XAI for science, and specify how to interface the machine view to domain experts and scientists. We note that the important assumption on the model being causal made at the beginning of this section must hold beckers2022causal ; carloni2023role . By having these pillars in place, scientists may grasp divergent views, increase their understanding of a given problem, and possibly fill knowledge gaps.
Figure 2: XAI for Science workflow. A predictive AI outperforms human expertise (grey box). Interpretability is then applied to generate the machine view (red boxes) that meet accuracy, reproducibility, and understandability (ARU) criteria. These interpretability results are then analyzed by human domain experts or generative AI models (green box) to derive interpretability-guided explanations (IGEs), which can either align with existing knowledge or lead to new scientific discoveries (in yellow). We depict two examples, one in the context of medicine (ECG data) and one in the context of climate science (weather data).
Let us now take two practical examples to outline the workflow, as depicted in Figure 2: (i) ECG time series signals for identifying a possible underling pathology of a given patient, and (ii) spatio-temporal weather data to forecast if an extreme event is or not occurring tomorrow. For both datasets, we use Transformer neural networks, that we deem casual for the sake of showcasing the possible workflow, and they we setup (predictive) classification tasks: classify pathology A or B for a given patient, and classify if an extreme event is happening tomorrow or not.
The first step we perform is to obtain classification scores, and evaluate whether these score are competitive, and the predictive AI outperforms human knowledge and predictive abilities. We then apply post-hoc interpretability, and assess that the relevance maps produced satisfy the ARU (accuracy, reproducibility, and understandability) requirements we just set. We then pass these interpretability results to human domain experts (cardiologists and meteorologists, respectively), and/or generative AI ‘scientists’ (large language models), and derive interpretability-guided explanations (also referred to as IGEs). These interpretability-guided explanations or IGEs can then be compared to existing human knowledge on the subject and they may generate nothing new, or could yield knowledge discovery.
From the two examples just introduced, we shall note some limitations and potential issues of the proposed methodology, that require further AI developments – these are discussed next.
Limitations and future outlook
XAI for Science faces some obvious limitations, related to both the causality assumption, as well as to the ARU requirements we set. In addition, translating the explanations provided by domain experts, and potentially by large language models, into actionable findings is still a nascent field that demands significant developments over the next few years.
Let us start from the assumption of causality. The patterns and relationships identified by the machine may be just spurious correlations, without any causality, and may fool scientists towards chasing wrong answers to the ‘why’ question. Understanding whether causality is appropriate to the problem being investigated, and assessing whether the machine view is causal is a key aspect that scientists should take into account in answering ‘why’ the machine deemed important certain data. Indeed, several works are now exploring the intersection between XAI and causality carloni2023role , two fields that in the computer science community have drifted apart over the years beckers2022causal , but that should be kept together, especially in the natural and applied sciences. To this end, some promising works are emerging – see e.g., scholkopf2022causality ; janzing2020feature ; xu2020causality . In addition, Peters et al. book peters2017elements provides an introduction to causality for learning algorithms, while Pearl’s book offers an accessible roadmap and perspective on causal inference from both a statistical and a philosophical perspective pearl2018book .
In terms of the accuracy requirement, to obtain what we called the machine view there exist several methods, grouped into two different categories, namely post-hoc and ante-hoc interpretability – see e.g., graziani2023global for a comprehensive review and taxonomy of interpretable AI. Post-hoc interpretability suffers from inaccuracy, aspect that may render futile the effort by domain experts to explain why certain data was used, if the data pinpointed is wrong. To this end, accuracy evaluation methods are as important, see e.g., turbe2023evaluation . Intrinsic interpretability may address the accuracy issues that plague post-hoc interpretability, yet they may be model-specific, human-constrained, and unable to provide machine views for complex deployed models that may not be easily accessible. Yet, several promising works are being carried out in the context of intrinsic interpretability – see e.g., alvarez2018towards ; zhang2022protgnn ; turbe2024protos . Another recent promising area is (intrinsic) compositional interpretability tull2024towards , that may help tie together different intrinsic interpretability viewpoints under a unified framework.
In terms of reproducibility, we should be able to retrieve the same machine view repeatedly, for a given task and set of samples. In this context, some issues derive from the possible non-uniqueness of the machine view that interpretability methods may provide (especially post-hoc interpretability methods). This can hinder our understanding of why those machine views were used. To this end, answering the question ‘how’ was the machine view (i.e., the data deemed important) used by the machine can provide a more useful insight, as well pointed out by Rudin rudin2019stop .
In terms of understandability, the machine view should allow scientists and domain expert to tap into their body of knowledge, to allow for comparing machine and human views on a certain phenomenon of interest. This is partly connected to what discussed by Freiesleben et al. freiesleben2022scientific . More specifically, the model should be able to address a question understandable to scientists, with a machine view that can also be understood by scientists within their framework of domain-specific knowledge. The choice of data and data configuration is therefore critical, and frequently arbitrary, leaving a degree of subjectivity and arbitrariness to the problem. For instance, if we consider the growing field of domain adaptability (i.e., using an AI model for a different domain than the one it was trained on) long2015learning , the machine view may not be understandable to scientists. These issues are less pronounced in the context of fine-tuning a certain model, as the pre-training is usually accomplished within the same domain as the fine-tuning, hence understandable to domain experts working in the specific field.
Finally, how do we make potentially useful interpretability-guided explanations (IGEs) provided by domain experts actionable – in other words, how do we use them in practice for knowledge discovery? A key field that has consistently driven human innovations is our ability to synthesize mathematical models of the real world galileo . However, from IGEs to mathematical models there is still a significant gap, with symbolic regression (e.g., angelis2023artificial ) possibly constituting feasible pathway.
Without a mathematical model, we can still grasp useful and actionable insights. For instance, Strogatz, in its inspiring New York Times editorial covering the winning of AlphaZero against Stockfish strogatz2018one states: “What is frustrating about machine learning, however, is that the algorithms can’t articulate what they’re thinking. We don’t know why they work, so we don’t know if they can be trusted. AlphaZero gives every appearance of having discovered some important principles about chess, but it can’t share that understanding with us.”. He additionally cites Garry Kasparov (the former world chess champion) that stated: “we would say that its [AlphaZero] style reflects the truth. This superior understanding allowed it to outclass the world’s top traditional program despite calculating far fewer positions per second.” kasparov2018chess . A few years later from that editorial, researchers showed that Go players are improving their game looking at how AlphaGo, DeepMind’s Go machine is playing, and providing them with a divergent view shin2023superhuman .
While science is a rather different application than Go, we believe that leveraging explanation for scientific inquiry that may ultimately lead to knowledge discovery may not be a foolish path. XAI is a possible path forward to keep complex human reasoning in the loop, possibly complemented by AI reasoning in large language models, while bridging current AI power in discovering patterns and relationships in data. XAI may also alleviate some of the risks that we may face when using AI for scientific discovery, that we share with Messeri and Crockett messeri2024artificial .
Acknowledgments
We want to thank my two PhD students, Jiawen Wei and Chenyu Dong for many of the results presented in this work. We thank all members of my group at NUS, MathEXLab, for their dedication to research that made this work possible. We also thank all the colleagues that provided constructive criticism, from several different research fields and perspectives, that helped improve this manuscript significantly. We acknowledge funding from MOE Tier 2 grant 22-5191-A0001-0 ‘Prediction-to-Mitigation with Digital Twins of the Earth’s Weather’ and MOE Tier 1 grant 22-4900-A0001-0 ‘Discipline-Informed Neural Networks for Interpretable Time-Series Discovery’.